### v2

@@ Abstract
We provide Dutch-book arguments against two forms of misspecified Bayesian learning. An agent progressively learns about a state and is offered a bet after every new discovery. We say the agent is Dutch-booked when they are willing to accept all bets, but their payoff is negative under each state either ex-post, or in expectation given the objective conditional probabilities of the discoveries (i.e., the correct data-generating process, DGP). Respectively, an agent cannot be Dutch-booked if and only if they update their beliefs with Bayes rule either from the previous belief, even using misspecified likelihood functions, or from one lexicographic prior, using the correct data-generating process. Under a large population interpretation of the DGP, this means that a population can suffer aggregate losses under all states when different individuals update their beliefs from different (lexicographic) priors, or using misspecified likelihoods. Thus, the Dutch-book argument offers a general characterization of the perils of misspecification.
@@ 1 Introduction
The classical Dutch-book argument (Ramsey 1931, De Finetti 1937) is a cornerstone of probability theory. It says that an agent (he) may accept a set of bets that return an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief. Later (Lehman 1955, Kemeny 1955), the converse was also proven: an agent cannot be Dutch-booked if he evaluates all bets with the same probabilistic belief about the state.
@@
In this Dutch-book argument there is no time dimension. But the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet after every piece of news. In this paper, we consider the following learning environment. There are a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices. The fundamental states represent different (probabilistic) DGPs for contingencies. Thus, the contingencies have objective probabilities conditional on the fundamental states. Then, a question arises: Should the beliefs at counterfactual contingencies be somehow coherent to guarantee sound decision-making? We answer this question with a two-fold Dutch-book argument. [^1] Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space (see Rescorla, 2022 for a recent survey). Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning. Rescorla (2022) extends the argument to the presence of subjectively impossible states. While in the partitional framework each state can only give rise to one learning path, and thus counterfactual contingencies can only be consistent with different states, in the framework of this paper a state can result into different paths, and thus be consistent with different counterfactual contingencies.
@@
We call deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is non-positive under all fundamental states and paths, and negative under some fundamental state and path. More generally, we call Dutch book a system of bets that yields a non-positive expected payoff under every fundamental state, and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration. An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (resp. deterministic) Dutch book. [^2] We use the following tie-breaking rule: the agent does not accept bets with zero subjective expected payoff.
In other words, a deterministic Dutch book guarantees that under every possible DGP the agents surely receives a negative payoff, and possibly gets a strictly negative payoff. Instead, a Dutch book requires that the agent’s expected payoff given the probabilities over contingencies induced by every DGP is nonpositive, and strictly negative for some DGP.
The notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning. With this, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates beliefs with Bayes rule whenever possible. Forward-consistency does not require to use the correct conditional probabilities of the contingencies, so the agent cannot be deterministically Dutch-booked even if they are misspecified about the relative probabilities in the likelihood function, i.e., they may update on the basis of a wrong likelihood function that links the fundamental states and the contingencies. It only rules out the more radical form of misspecification where a state is subjectively believed possible given some contingency that is objectively impossible under that state. Moreover, forward-consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each current belief is a convex combination of the beliefs at the next contingencies. Thus, Bayes-plausibility is the coherence rule among beliefs at counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked.
@@
The necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework (see Rescorla, 2022 for the most general case), because of the multiplicity of learning paths that are compatible with the same state. Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the next contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework. [^3] Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on one dimension of the “grand state space” that also includes the learning paths. The solution to this problem will be an application of Farkas lemma.
Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, the general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state. Such weaker form of Dutch book is important for two orders of reason.
First, a population of agents making choices in different contingencies will realize the aggregate payoff under the true state. So, if the population can be Dutch-booked in objective expected terms, their combined choices will select sure aggregate losses and forfeit sure aggregate wins, suggesting an evolutionary disadvantage. Alternatively, a single agent may perish in the long run after being exposed to i.i.d. occurrences of the same dynamic decision problem.
Second, the currently accepted interpretation of Dutch-book arguments is not the literal one, where a bookmaker can truly design a bet for each contingency, but rather that the agent suffers from some form of irrationality, which the Dutch-book argument brings to light.
We illustrate how general Dutch books work with an example.
Example 1
Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding. There are three possible locations: main square, marina, and central park. There are three bus lines on his street: the red line goes to the square, the blue line to the marina, and the green line to the park. The buses always arrive in this order, at constant time intervals. Larry is late, so his only chance to see the fireworks is taking the first bus that arrives, regardless of its color, and hoping it goes to the right location. Suppose the blue line comes. When Larry arrives at the marina, he discovers whether the fireworks are there or not. If they are not, Larry regrets not moving to the bus stop a little earlier, which meant catching the red line: “I bet the fireworks are at the square!” Overhearing this, a bookmaker approaches Larry and proposes a bet on the fireworks’ location. Larry accepts a bet that yields $9$ if the fireworks are at the square and $-10$ if the fireworks are at the park. By the same token, at the park, he would accept a bet that yields $9$ for marina and $-10$ for square, and at the square, he would accept a bet that yields $9$ for park and $-10$ for marina. Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior.
In the example, the bus schedule determines the objective conditional probabilities of the three contingencies of interest. Instead, there are no objective probabilities of the fireworks’ location, and Larry does not have a subjective prior either; after all, he cannot decide which bus to take, so he does not need one. However, to deal with the bookmaker, he needs to form a belief after discovering that the fireworks are not in the location he arrived to. What’s wrong with the beliefs he forms? The location he deems more likely is different in each of the three contingencies. This form of incoherence among beliefs can arise precisely from the lack of a prior. Thus, the Dutch-book argument provides a motivation for having a prior, even when there is no decision to make before the arrival of information. More generally, the argument provides a justification for having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one, but just reason in a coherent way after different discoveries. In the population interpretation, having a prior translates into the agents having the same prior. Thus, the Dutch book arguments suggests a disadvantage for “divided” populations, in which different agents have different (prior) beliefs, in that the agents with different priors can be exploited in different situation. Such divided population naturally emerges in a repeated setting if the agents do not share their experiences, thus suggesting an important role for the aggregation of feedback.
Another possible reason for Larry’s incoherent beliefs is that he (has a prior, but) perceives an inexistent correlation between bus arrivals and location of the fireworks. This is a classic form of model misspecification about the likelihood function connecting fundamental states and contingencies. In fact, we show that any form of model misspecification about the likelihood function allows Dutch-booking. This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states. The Dutch-book argument offers a precise sense in which model misspecification is bad for decision-making.
In this simple example, having one prior and updating it by conditioning suffices to avoid Dutch-booking. But this is not true in general. We determine the coherence rule among beliefs that is necessary and sufficient to avoid Dutch-booking in objective expected terms. The coherence rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states. In the example, if Larry deems square twice more likely than park when he is at the marina and marina twice more likely than square when he is at the park, to satisfy the coherence rule he should deem marina four times more likely than park when he is at the square.
In turn, we show that this coherence rule characterizes a Completely Consistent Belief System: an array of beliefs (one for each contingency) that can be derived from a Lexicographic Conditional Probability System (Blume et al., 1991), a list of theories in a priority order. Specifically, each belief is derived with Bayes rule from the first theory that is consistent with the contingency. So, the LCPS plays the role of lexicographic prior, while the CCBS specifies the posterior beliefs in different contingencies. The fact that a completely consistent belief rather than simply conditioning from a unique prior is needed will be illustrated in Section 6 with an example of an agent who is subject to correlation neglect.
@@
In the context of sequential games, Siniscalchi (2022) introduces the notion of structural rationality, which is based on the use of CCBSs over the opponents’ strategies. [^4] Interestingly, structural rationality is motivated by the inability of sequential rationality to capture coherent conditional preferences that can be derived by updating a prior preference relation. CCBSs guarantee that all conditional preferences can be derived from a prior lexicographic preference relation. In other words, they guarantee a stronger form of Bayesianism which might be called “Bayesian updating of preferences” compared to Bayesian updating of beliefs. In Siniscalchi (2022), as well as Battigalli et al. (2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System: an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability. [^5] Conditional probability systems were first introduced by Renyi (1955). CCPSs were introduced in the analysis of games by Myerson (1986). In turn, a CCPS can be derived from a LCPS by conditioning. In a sequential game, a player progressively rules out opponents’ strategies by observing their moves, that is, learning occurs in the partitional fashion. So, Bayes rule boils down to conditioning, and hence the definition of CCBS in Siniscalchi (2022) and Battigalli et al. (2023) is a special case of the definition of this paper. Thus, the Dutch-book argument for complete consistency implies that the standard notion of conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms. The working paper of Siniscalchi (2022) also characterizes CCBSs with a generalization of the chain rule that applies between non-nested conditional events. In the Appendix, we provide a direct proof of the equivalence between the generalized chain rule and our coherence rule in sequential games.
@@
Pettigrew (2023) considers an agent with a credence at time $t$ and a vector of possible credences at $t^{\prime}>t$, and shows that the agent can be Dutch-booked if and only if his credence at $t$ is not in the convex hull of the credences at $t^{\prime}$ (which he calls “General Reflection Principle”). This result yields our equivalence between Bayes-plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state. [^6] When some discoveries objectively rule out some states, statement II.a of Theorem 1 in Pettigrew (2023) fails. From his theorem, Pettigrew also derives an argument for conditioning in the partitional framework. In our richer framework, we can derive our more general argument for forward consistency, uncovering the limited possibility of achieving a deterministic Dutch-booking against likelihood misspecification.
@@
Chen et al. (2015) show that the posteriors of different agents, who update a common prior based on different discoveries, are common knowledge if and only if they cannot be Dutch-booked in aggregate. [^7] See also Lehrer and Samet (2011) for an alternate characterization of “agreeing to agree” between only two agents. Their notion of Dutch book only takes the realized discoveries and sums up the payouts under each state. This amounts to assuming that the bookmaker can choose the bets based on the entire set of realized discoveries. So, when the agents have different (thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each single agent, who is unaware of the other agents’ discoveries. In other words, the bookmaker can make profit from arbitrage. This is ruled out by our notion of Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents. This reconciles the possibility of Dutch-booking Bayesian agents in Chen et al. (2015) with the impossibility results of this paper.
@@
There are several related contribution in the literature on misspecified Bayesian learning, who is excellently surveyed in Bohren and Hauser (2024a). Lanzani (2025) characterizes a behavior rule that prevents misspecification to damage the agent beyond his maxmin payoff preserving consistency when correctly specified. Bohren and Hauser (2024b) characterizes when a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as the consequence of Bayesian updating within a misspecified model. Molavi (2024) shows that beliefs are consistent with Bayesian updating (from a possibly misspecified model) if and only if the mean posterior is absolutely continuous with respect to the prior. This finding generalizes the earlier Shmaya and Yariv (2016) to the case in which the subjective beliefs of the decision maker may not have the same support of the true data-generating process. Cerreia-Vioglio et al. (2025) show that in coordination games misspecification can be extremely damaging even for misspecification concerned agents. Compared to this literature, we provide a novel reason for why some misspecification should or should not be expected to persist based on their susceptibility to Dutch books.
@@
Finally, our results are similar in spirit to those in the no-trade literature, the closest being in Morris (1994). Morris considers agents who trade claims over commodities that are contingent on a payoff-relevant state and on the realization of a signal. In the case of two traders, a single money commodity and risk neutrality, Morris’ model coincides with ours given a tree of depth $1$. Assuming that the traders have no incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals. Our main theorem further shows (in a more general learning framework) that such trade can take a very specific form, whereby one trader has a positive expected payoff conditional on each state, and the other has a positive expected payoff conditional on each signal. If the first trader (the bookmaker) is using the correct likelihood function, then the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked. Conversely, we show that the existence of such trade guarantees that the disagreement between the two traders must be about the likelihood function, not (just) the prior.
The paper is organized as follows. Section 1 introduces the framework. Section 2 illustrates belief systems. Section 3 introduces the novel notions of Dutch book. Section 4 states and proves the Dutch-book theorem for forward-consistency. Section 5 is devoted to the Dutch-book theorem for complete consistency. In the Appendix, we formalize the special case of sequential games and the connection with Siniscalchi (2022).

### v3

@@ Abstract
We provide Dutch-book arguments against misspecified Bayesian learning. An agent progressively learns about a state and is offered a bet after every discovery. We say the agent is deterministically Dutch-booked when they would accept all bets, but their payoff is ex-post negative under each state. More generally, we say that the agent is Dutch-booked when they would accept all bets, but their expected payoff under each fundamental state is negative. With this, the agent cannot be deterministically Dutch-booked if and only if they update their beliefs using Bayes’ rule, even with misspecified likelihoods. The agent cannot be Dutch booked if and only if they update their beliefs using Bayes’ rule with a lexicographic prior and using the correct data-generating process. We show that offers of financial instruments and behavior in Monty Hall problems can be viewed as Dutch books that extract a sure expected gain from a misspecified population.
@@ 1 Introduction
A growing literature (surveyed in Bohren and Hauser, 2025) documents the consequences of updating beliefs using a misspecified model of the world. Model misspecification leads entirely rational (i.e., Bayesian) agents astray by inducing them to misinterpret the data and compute incorrect posteriors. Is there a general sense in which model misspecification is harmful for decision-making? We provide an answer to this question through a two-fold Dutch-book argument.
@@
The classical Dutch-book argument (Ramsey 1931, De Finetti 1937) is a cornerstone of probability theory. It says that an agent (he) may accept a set of bets that returns an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief. Later (Lehman 1955, Kemeny 1955), the converse was also proven: an agent cannot be Dutch-booked if he evaluates all bets with the same probabilistic belief about the state. In the classical Dutch-book argument, there is no time dimension. However, the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet conditional on every piece of news. Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space on which he bets. [^1] An exception is Pettigrew (2023): see the literature review. Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning. [^2] Rescorla (2022) surveys Dutch-book arguments and extends the argument for conditioning to the presence of subjectively impossible states.
However, in many problems of dynamic choice, the decision-relevant state does not include the observed events; hence, learning is best represented as the acquisition of noisy signals. That is the approach that is followed in the entire economics literature on misspecified learning. In particular, this literature employs the following learning environment, which we also adopt here. There is a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices. The fundamental states represent different (probabilistic) DGPs for contingencies. Thus, the contingencies have objective probabilities conditional on the fundamental states. Then, the question arises: what if the agent is (Bayesian but) misspecified about these probabilities?
@@
We call a deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is nonpositive under all fundamental states and paths, and negative under some fundamental state and path. [^3] Thus, the notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning. More generally, we call a system of bets a Dutch book if it yields a nonpositive expected payoff under every fundamental state and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration. An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (deterministic) Dutch book. [^4] We use the following tie-breaking rule: the agent does not accept bets with zero subjective expected payoff. In other words, a deterministic Dutch book guarantees that, under every possible DGP, the agent surely receives a nonpositive payoff and possibly a strictly negative payoff. Instead, a Dutch book requires that the agent’s expected payoff, given the probabilities over contingencies induced by every DGP, is nonpositive and strictly negative for some DGP.
@@
First, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates his beliefs using Bayes rule whenever possible. Since forward-consistency does not require the use of the correct conditional probabilities of the contingencies, agents that are Bayesian but misspecified about the relative probabilities in the likelihood function (i.e., they may apply Bayes rule to the wrong likelihood function linking states and the contingencies) are immune to deterministic Dutch books. However, an agent with the more radical form of misspecification, in which a state is subjectively believed possible given a contingency objectively impossible under that state, can be deterministically Dutch-booked. Moreover, forward consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each belief is a convex combination of the beliefs that immediately follow. Thus, Bayes-plausibility is the coherence rule among beliefs in counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked. [^5] Note that the necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework, because of the multiplicity of learning paths that are consistent with the same state. Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the following contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework. Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on only one dimension of the “grand state space,” which also includes the learning paths. The solution to this problem will be an application of Farkas lemma.
Second, and most importantly, we show that a Bayesian agent cannot be Dutch-booked (“in objective expected terms”) if and only if he updates beliefs from a lexicographic prior with the correct likelihood function. This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states. Our Dutch-book argument offers a precise sense in which model misspecification is harmful for decision-making: a misspecified agent could accept sure expected losses and forfeit sure expected wins. The argument also offers a motivation for having a prior, even when there is no decision to take before the arrival of information. The use of a lexicographic prior (in place of free belief revision after each surprise) is crucial to guarantee the coherence of beliefs across counterfactual contingencies, thereby shielding the agent from a Dutch book. We call a belief system completely consistent (henceforth CCBS) when it is derived from a subjective lexicographic prior with the correct likelihood function. More generally, the argument justifies having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one; rather, the agent may reason coherently after different discoveries. We call complete coherence the rule of coherence between beliefs in counterfactual contingencies that shelters the agent from Dutch-booking. The rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states.
Importantly, we also show that the possibility of Dutch-booking a misspecified agent does not rely on the bookmaker knowing the true data generating process perfectly. As long as a well defined measure of discrepancy between the agent and bookmaker beliefs is sufficiently large compared to the bookmaker misspecification, a Dutch book can be constructed.
Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, our general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state. Besides being a useful theoretical tool to uncover a form of “irrationality”, this weaker form of Dutch book could also be of applied relevance, even though a single agent will only find himself in one of many counterfactual contingencies. First, a large population of agents making choices in different contingencies will realize the aggregate payoff under the actual state. So, our Dutch-book argument suggests a disadvantage for “divided” populations, in which different agents hold different (prior) beliefs and can thus be exploited in different situations. Such a divided population naturally emerges in a repeated setting when agents do not share their experiences, suggesting an important role for aggregating feedback. Second, a single agent may perish in the long run after being exposed to i.i.d. occurrences of the same dynamic decision problem. Third, an agent may be offered the whole set of conditional bets at the ex-ante stage, as in a speculative-trade context (see the related literature below). Traditional Dutch-book arguments are often criticized for the possibility that the agent would realize being Dutch-booked when all bets are offered simultaneously. However, in our case, a Dutch-booked agent will still compute a positive expected payoff under some state by using a misspecified likelihood function. [^6] Dutch-book arguments have also been criticized for the need of a bookmaker who knows the agent’s beliefs. We disagree with this critique, as a bookmaker may find a Dutch book that the agent wants to accept by trial and error. In a financial market, the agent may also self-pick a set of assets that jointly constitute a Dutch book, without the presence of a strategic bookmaker.
We then apply our theorems to one commonly studied form of misspecification, correlation neglect. Since agents with this bias update beliefs using Bayes rule, our first result implies that they are shielded from deterministic Dutch books. However, they use an incorrect likelihood that neglects the correlation between contingencies, thereby exposing them to general Dutch books, consistent with our second result. In particular, we illustrate the nature of those Dutch books in a venture capital setting and in a variation of the Monty Hall game, with secretly assigned doors. In the former application, the Dutch book consists of the venture capitalist offering different combinations of convertible debt and ownership purchases depending on whether initial milestones are achieved. In the second case, the Dutch book offers terms that seem inviting if the remaining doors were equally likely, but that induce sure expected loss under the mechanism used to determine what door is open initially.
@@
In the context of sequential games, Siniscalchi (2022) introduces the notion of structural rationality, which is based on the use of CCBSs over the opponents’ strategies. [^7] Interestingly, structural rationality is motivated by the inability of sequential rationality to capture coherent conditional preferences that can be derived by updating a prior preference relation. CCBSs guarantee that all conditional preferences can be derived from a prior lexicographic preference relation. In other words, they guarantee a stronger form of Bayesianism, which might be called “Bayesian updating of preferences” rather than “Bayesian updating of beliefs.” In Siniscalchi (2022), as well as Battigalli et al. (2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System (CCPS): an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability. [^8] Conditional probability systems were first introduced by Renyi (1955). CCPSs were introduced in the analysis of games by Myerson (1986). In turn, a CCPS can be derived from a lexicographic probability system by conditioning. In a sequential game, a player progressively rules out opponents’ strategies by observing their moves; that is, learning occurs in a partitional fashion. So, Bayes rule boils down to conditioning, and hence the definition of CCBS in Siniscalchi (2022) and Battigalli et al. (2023) is a special case of the definition of this paper. Thus, the Dutch-book argument for complete consistency implies that the standard notion of a conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms. The working paper of Siniscalchi (2022) also characterizes CCBSs by generalizing the chain rule to apply to non-nested conditional events. In the Appendix, we provide a direct proof of the equivalence between the generalized chain rule and our coherence rule in sequential games.
@@
Pettigrew (2023) considers an agent with a credence at time $t$ and a vector of possible credences at $t^{\prime}>t$, and shows that the agent can be Dutch-booked if and only if his credence at $t$ is not in the convex hull of the credences at $t^{\prime}$ (which he calls “General Reflection Principle”). This result yields our equivalence between Bayes plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state. [^9] When some discoveries objectively rule out some states, statement II.a of Theorem 1 in Pettigrew (2023) fails. From his theorem, Pettigrew also derives an argument for conditioning in the partitional framework. In our richer framework, we can derive our more general argument for forward consistency, establishing the impossibility of achieving a deterministic Dutch-booking against likelihood misspecification.
@@
Chen et al. (2015) show that the posteriors of different agents, who update a common prior based on different discoveries, are common knowledge if and only if they cannot be Dutch-booked in aggregate. [^10] See also Lehrer and Samet (2011) for an alternate characterization of “agreeing to agree” between only two agents. Their notion of a Dutch book considers only realized discoveries and sums up the payouts under each state. This amounts to assuming that the bookmaker can choose the bets based on the entire set of realized discoveries. So, when the agents have different (and thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each agent, who is unaware of the other agents’ discoveries. In other words, the bookmaker can make a profit from arbitrage. This is ruled out by our notion of a Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents. This reconciles the possibility of Dutch-booking Bayesian agents in Chen et al. (2015) with the impossibility results of this paper.
Molavi (2025) shows that beliefs are consistent with Bayesian updating if and only if the mean posterior is absolutely continuous with respect to the prior. This finding generalizes the earlier work of Shmaya and Yariv (2016) to the case in which the decision maker’s subjective beliefs may not have the same support as the true data-generating process. Fudenberg and Lanzani (2026) characterize when beliefs are consistent with Bayesian updating from a conditional i.i.d. model.
@@
There are several related contributions in the literature on misspecified Bayesian learning, which is excellently surveyed in Bohren and Hauser (2025) and Esponda and Pouzo (2026). Here we discuss the part of this literature concerned with measuring the loss induced by misspecification and how to avoid these losses. Lanzani (2025) characterizes a behavior rule that prevents misspecification from damaging the agent beyond his maxmin payoff, preserving consistency when correctly specified. Frick et al. (2024) provide a rank of forms of misspecification on the basis of the damages they induce, and Gossner and Steiner (2018) provides a measure of cost of misperceiving the payoff function. Bohren and Hauser (2024) characterize the conditions under which a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as a consequence of Bayesian updating within a misspecified model. Cerreia-Vioglio et al. (2025) show that misspecification in coordination games can be highly damaging, even for agents concerned about it. Compared to this literature, we provide a novel reason for why some misspecification should or should not be expected to persist based on their susceptibility to Dutch books. Moreover, as our application to correlation neglect clearly shows, the form of misspecification that are susceptible to Dutch book significantly differ from the one that have been deemed unstable due to light-bulb realizations or evolutionary pressure (see, Gagnon-Bartsch et al. 2023, Fudenberg and Lanzani 2023, Ba 2026, He and Libgober 2025).
@@
Finally, our results are similar in spirit to those in the no-trade literature, the closest being in Morris (1994). Morris considers agents who trade claims over commodities that are contingent on a payoff-relevant state and on the realization of a signal. In the case of two traders, a single money commodity, and risk neutrality, Morris’ model coincides with ours with a tree of depth $1$. Assuming that the traders have no ex-ante incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals. Our main theorem further shows (in a more general learning framework) that such trade can take a very specific form, whereby one trader has a positive expected payoff conditional on each state, and the other has a positive expected payoff conditional on each signal. Suppose the first trader (the bookmaker) is using the correct likelihood function. In that case, the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked. Conversely, we show that the existence of such trade guarantees that the disagreement between the two traders must be about the likelihood function, not (just) the prior.